answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of getting one head when flipping a coin three times?

The probability is 3/8.

The probability is 3/8.

The probability is 3/8.

The probability is 3/8.

What is probability of getting a 3 in a deck of cards?

For an ordinary deck of cards, the probability is 1. All decks of playing cards contain 3s.

For an ordinary deck of cards, the probability is 1. All decks of playing cards contain 3s.

For an ordinary deck of cards, the probability is 1. All decks of playing cards contain 3s.

For an ordinary deck of cards, the probability is 1. All decks of playing cards contain 3s.

What is the probability that the number rolled will be greater that 4?

The probability that the number rolled, on a fair, six sided die, will be greater than 4 is 1/3.

The probability that the number rolled, on a fair, six sided die, will be greater than 4 is 1/3.

The probability that the number rolled, on a fair, six sided die, will be greater than 4 is 1/3.

The probability that the number rolled, on a fair, six sided die, will be greater than 4 is 1/3.

Why is the Gaussian distribution referred to as normal distribution?

Because many naturally occurring variables were approximately distributed according to a Normal bell shaped curve.

Because many naturally occurring variables were approximately distributed according to a Normal bell shaped curve.

Because many naturally occurring variables were approximately distributed according to a Normal bell shaped curve.

Because many naturally occurring variables were approximately distributed according to a Normal bell shaped curve.

What is the probability of a coin landing heads up on two consecutive flips?

In an infinite series of flips it is 1 = a certainty.

In only two flips it is 1/4.

In an infinite series of flips it is 1 = a certainty.

In only two flips it is 1/4.

In an infinite series of flips it is 1 = a certainty.

In only two flips it is 1/4.

In an infinite series of flips it is 1 = a certainty.

In only two flips it is 1/4.

How many ways can the letters in the word country be arranged?

Using all letters? You have seven different letters, so you can rearrange them in 7! (read: seven factorial) different ways, that is, 1 x 2 x 3 x 4 x 5 x 6 x 7.

What is the probability of getting exactly three heads when you flip a coin four times?

If it is a fair coin, the probability is 1/4.

If it is a fair coin, the probability is 1/4.

If it is a fair coin, the probability is 1/4.

If it is a fair coin, the probability is 1/4.

What is the probability of rolling a 1 2 or 3 on the first roll of a 1-6 number cube and rolling a 4 5 or 6 on the second roll of the same cube?

The probability of (1 or 2 or 3) on the first (or any) roll is 1/2 = 50% .

The probability of (4 or 5 or 6) on the second (or any) roll is 1/2 = 50% .

The probability of exactly that result on two rolls is (1/2 x 1/2)= (50% x 50%) = 1/4 = 25% .

Why do you think is probability used as basis for inferential statistics?

Inferential statistics are used in situations where it can be assumed that random behaviour(s), subject to the mathematical laws of probability, must be taken into account.

What is the probability of rolling divisors of 12 on both dice?

Of all the six numbers on a cube, only the '5' is not a factor of 12.

Probability of rolling a factor of 12 on the first cube = 5/6 .

Probability of rolling a factor of 12 on the second cube = 5/6 .

Probability of rolling a factor of 12 on both cubes = (5/6 x 5/6 ) = 25/36 = 694/9 %.

What is the intersection ofthe assessed probability and severity of a hazard called in the crm process?

It is risk assessment.

It is risk assessment.

It is risk assessment.

It is risk assessment.

How many different ways can six people - ABCDEF - sit in a row at a movie theater if A and B will not sit next to each other?

There are 480 ways.

The simplest way to get the answer is to find the total number of ways that they can sit and, from that, subtract the number in which A and B are next to one another.

Total number of ways = 6! = 6*5*4*3*2*1 = 720.

Now suppose A and B are next to one another. Treat AB as a single unit, and then the total number of ways is 5! = 5*4*3*2*1 = 120. But A and B can be next to one another in two ways: AB and BA. So, the total number of ways with A and B next to one another is 2*120 = 240.

That leaves 720 - 240 = 480 ways in which the six can sit in a row with A and B not together.

What are the odds in favor of rolling a 7 or an 11 or doubles twice in a row with a pair of standard dice?

The odds of rolling a 7 are 1/6. The odds of rolling two in a row are 1/36.

The odds of rolling an 11 are 1/18. The odds of rolling two in a row are 1/324.

The odds of rolling doubles are 1/6. The odds of rolling double twice in a row are 1/36.

What is the probability that Jen will roll a 3 on a number cube and flip tails on a coin?

There is a 1/6 chance for the 3 and a 1/2 chance for the coin. Multiply them to get a 1/12 chance for both. ~.0833%

What is the probability of obtaining at least one head when a coin is flipped four times?

The probability of obtaining 4 tails when a coin is flipped 4 times is:

P(4T) = (1/2)4 = 1/16 = 0.0625

Then, the probability of obtaining at least 1 head when a coin is flipped 4 times is:

P(at least 1 head) = 1 - 1/16 = 15/16 = 0.9375

What is -1.33 standard deviations in percentage?

Assuming a normal distribution, Pr { X < -1.33 } ~= 0.091759135650280765

or about 9.18 %

What is the probability of selecting a king or queen?

It depends on the context which is not specified. If picking a topic for a history project on the monarchy of some country it would be 1!

If it concerns selecting one card, at random, from a standard deck of playing cards, the answer is 8/52 = 2/13.

What is discrete probability distribution?

It is a function that gives the probabilities associated with the discrete number of values that a random variable can take.

What happen to confidence interval if increase sample size and population standard deviation simultanesous?

The increase in sample size will reduce the confidence interval.

The increase in standard deviation will increase the confidence interval.

The confidence interval is not based on a linear function so the overall effect will require some calculations based on the levels before and after these changes.

It would depend on the relative rates at which the change in sample size and change in standard deviation occurred. If the sample size increased more quickly than then standard deviation, in some sense, then the size of the confidence interval would decrease. Conversely, if the standard deviation increased more quickly than the sample size, in some sense, then the size of the confidence interval would increase.

Trending Questions
How many queens ace or kings are there in a deck of cards? What is the probability of rolling a composite number with one roll of a number cube? All the factors of 160? Is the gender of college students a discrete random variable a continuous random variable or not a random variable? Two cards are drawn from a standard deck of 52 cards without replacement Find the probability that both cards are aces? What are the chances of an acorn falling on your head? Does the Student t distribution takes different values for a different sample size? What is the probability of choosing a marble that is blue? When all outcomes of an experiment are equally likely? What is the probability of a square? How many equal groups of cards can be made from a standard deck of 52 cards? Suppose you have an extremely unfair coin the probability of a head is ¼ and the probability of a tail is ¾ If you toss the coin 32 times how many heads do you expect to see? How do you find the probability of a percent? What are the chances of winning the euromillions? Two coins are tossed and a tetrahedral die is thrownthe tetrahedral die has 4 faces 1234 the probablity that only one head and even number? What two interests does probable cause balance? What is the probability of getting and odd number when rolling as single 6 sided die? What is the most probable ion formed for Br? What is the probability of three digit numbers without repeated digits with numbers 0123456? Using excel including histogram info how do I get the probablilty of 7 questions 4 possible answers 1 being correct the probalility of out of 200 students getting the correct asnwer?